OpenStudy (anonymous):

Calculate dx/dy, simplify y = (x^9)(squareroot of x)

OpenStudy (anonymous):

\[y = x^9\sqrt{x}\]

OpenStudy (anonymous):

(9x^8)(x^(1/2)/2)

OpenStudy (anonymous):

so the derivative of \[\sqrt{x}\] = (x^(1/2/2)

OpenStudy (anonymous):

Product rule. I don't believe the equation above is correct @dolanoscopy

OpenStudy (anonymous):

well the derivative of \[\sqrt{x}\] is (1/2)x^(1/2) which simplifies to x^(1/2)/2

OpenStudy (anonymous):

The derivative of root x is .5 x^-(1/2), you subtract 1 from 1/2

OpenStudy (anonymous):

Ahh, forgot the negative. My bad.

OpenStudy (anonymous):

The complete derivative of the question asked would be... (9x^8)(rootx) + (.5x^(-.5))(x^9) found using the product rule.